OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x2 + 1, find f(g(3)).

OpenStudy (p0sitr0n):

g(3)=3^1+1=9+1=10 f(g(3))=2*10+3=20+3=23